Solve for x log 5 (5 1/x + 125) = log 5 6 + 1 + 1/2x

Solve for x


log5 (51/x + 125) = log5 6 + 1 + 1/2x

Grade:12th Pass

2 Answers

Shubham Rohilla
11 years ago

Log(51/x + 125) = log(6) + log(5) + log (51/2x)

                     = log(6.5.51/2x)

                     = log(30.51/2x)

cancelling log from both sides we get

 52/2x + 125 = 30.51/2x   

  now put 51/2x = y ,

   we get,

 

y+ 125 = 30y

y2 - 30y + 125 = 0

(y - 5)(y - 25) = 0

from here we get y = 5 or 25

 so x can be 1/2 or 1/4 .

king arthur
11 years ago

Dude, you cannot just cancel log.. Appy 5 power on both sides